# Universal contributions to charge fluctuations in spin chains at finite temperature

### 17 Jan 2024 v1

@@ VI Acknowledgements
Note added: During the finalization of the manuscript, we became aware of closely related works Zang et al. and Wang et al. , which study the universal feature in $\alpha(\theta)$. In particular, Zang et al. showed that the O(2) LSM anomaly leads to the cusp at $\theta=\pi$ in $\alpha(\theta)$.
@@ Appendix A Typology of ’t Hooft anomaly in (1+1)d
Next we consider the Levin-Gu model, introducing a $\theta=\pi$ defect Cheng and Seiberg (2023):
@@ Appendix C $G(\theta)$ of the Levin-Gu model
We will approximate the discrete product in $Z_{ss^{\prime}}$ by integrals. Define
$I_{1}(\phi,\beta)$ $=\int_{0}^{2\pi}\mathop{}\!\mathrm{d}x\,\ln\left(1+\,\mathrm{e}^{\mathrm{i}{\phi}}\,\mathrm{e}^{4\beta\cos x}\right),$ (104)
$I_{2}(\beta)$ $=\int_{0}^{2\pi}\ln\left\lvert 1-\,\mathrm{e}^{4\beta\cos x}\right\rvert\mathop{}\!\mathrm{d}x.$ (105)
A crucial property of $I_{1}(\phi,\beta)$ for our derivations is that for $-\pi<\phi<\pi$,
$\Im I_{1}(\phi,\beta)=\pi\phi.$ (106)
@@
$\sum_{j=\frac{s^{\prime}}{2}}^{N-1+\frac{s^{\prime}}{2}}\ln\left(\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}+1\right)=\frac{N}{2\pi}\int_{0}^{2\pi}\mathop{}\!\mathrm{d}x\,\ln(\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{\alpha\cos x}+1)+\mathcal{O}(N^{-1}).$ (107)
@@
$Z_{0s^{\prime}}(\phi,\beta)\approx\exp\left[\frac{N}{2\pi}I_{1}(\phi,\beta)\right]=\,\mathrm{e}^{\frac{1}{2}\mathrm{i}[\phi]_{\pi}N}\exp\left[\frac{N}{2\pi}\Re I_{1}(\phi,\beta)\right].$ (108)
@@
We also need the following inequalities between the integrals. First note that
@@
which implies that
$I_{1}(0,\beta)\geq\Re I_{1}\left(\phi,\beta\right).$ (111)
Similarly
$\,\mathrm{e}^{4\beta\cos x}+1>\left\lvert\,\mathrm{e}^{4\beta\cos x}-1\right\rvert\Longrightarrow I_{1}(0,\beta)>I_{2}(\beta).$ (112)
@@ C.1 N even
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}$ $\approx\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]+\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2}+\pi,\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (114)
$\approx\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (115)
$=\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right].$ (116)
@@
$Z_{0s^{\prime}}(\pm\pi/2,\beta)=Z_{1s^{\prime}}(\mp\pi/2,\beta)=\exp\left[\frac{N}{2\pi}I_{1}\left(\pm\frac{\pi}{2},\beta\right)\right].$ (117)
Crucially, the integrals $I_{1}(\pm\frac{\pi}{2},\beta)$ satisfy $\Re I_{1}(\pm\frac{\pi}{2},\beta)=I(\beta)$, where $I(\beta)$ is given explicitly by
$I(\beta)=\int_{0}^{2\pi}\ln\sqrt{1+\,\mathrm{e}^{8\beta\cos x}}\mathop{}\!\mathrm{d}x.$ (118)
Therefore we have
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}\approx 2\cos\left(\frac{\pi N}{4}\right)\exp\left[\frac{N}{2\pi}I(\beta)\right].$ (119)
@@
$\approx 4\exp\left[\frac{N}{2\pi}I_{2}(\beta)\right].$ (123)
@@
$G(\theta)\approx\frac{\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right]}{\exp\left[\frac{N}{2\pi}I_{1}(0,\beta)\right]+2(-1)^{\frac{N}{2}}\exp\left[\frac{N}{2\pi}I_{2}(\beta)\right]}\approx\,\mathrm{e}^{\frac{1}{4}\mathrm{i}\theta N}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right]},$ (124)
@@
$G(\pi)\approx\frac{2\cos\left(\frac{\pi N}{4}\right)\exp\left[\frac{N}{2\pi}I(\beta)\right]}{\exp\left[\frac{N}{2\pi}I_{1}(0,\beta)\right]+2(-1)^{\frac{N}{2}}\exp\left[\frac{N}{2\pi}I_{2}(\beta)\right]}\approx 2\cos\left(\frac{\pi N}{4}\right)\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-I(\beta)\right]}.$ (125)
@@ C.2 N odd
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}$ $\approx\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]-\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2}+\pi,\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (127)
$\approx\exp\left[\frac{N}{2\pi}I_{1}\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (128)
$=\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right],$ (129)
@@
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}\approx$ $\left[\exp\left(\frac{\mathrm{i}\pi N}{4}\right)-\exp\left(\frac{3\mathrm{i}\pi N}{4}\right)\right]\exp\left(\frac{N}{2\pi}I(\beta)\right)$ (130)
$=$ $2(-1)^{\frac{N-1}{2}}\sin\left(\frac{\pi N}{4}\right)\exp\left[\frac{N}{2\pi}I(\beta)\right].$ (131)
The large $N$ behavior of the partition function is still governed by $I_{1}(0,\beta)$ as $I_{1}(0,\beta)>I_{2}(\beta)$. Moreover, the two terms $Z_{10}(0,\beta)$ and $Z_{11}(0,\beta)$ have opposite sign for $N$ odd and they cancel each other. Write $N=4m+r$ where $r=1,3$, then $\sin\frac{\pi N}{4}=\sin(m\pi+\frac{r\pi}{4})=(-1)^{m}\sin\frac{r\pi}{4}=\frac{1}{\sqrt{2}}(-1)^{m}$. We conclude that for $0<\theta<\pi$,
$G(\theta)\approx\frac{\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right]}{\exp\left[\frac{N}{2\pi}I_{1}(0,\beta)\right]}=\,\mathrm{e}^{\frac{1}{4}\mathrm{i}\theta N}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-\Re I_{1}\left(\frac{\theta}{2},\beta\right)\right]},$ (132)
@@
$G(\pi)\approx\frac{(-1)^{m+\frac{r-1}{2}}\sqrt{2}\exp\left[\frac{N}{2\pi}I(\beta)\right]}{\exp\left[\frac{N}{2\pi}I_{1}(0,\beta)\right]}=(-1)^{m+\frac{r-1}{2}}\sqrt{2}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-I(\beta)\right]}.$ (133)
@@ D.1 Half filling
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}\approx\exp\left[\frac{N}{2\pi}I_{1}\left(\theta,\beta\right)\right]=\exp\left(\frac{1}{2}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I_{1}\left(\theta,\beta\right)\right].$ (135)
@@
We have proved that $\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}=0$ for $N$ odd, so we only need to consider $N$ even case. $Z_{00}(0,\beta)$ and $Z_{01}(0,\beta)$ can be approximated by Eq. (108), and they cancel each other. The other two terms are already considered in Appendix C, and we have $\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}=2(-1)^{\frac{N}{2}}\exp\left[\frac{N}{2\pi}I_{2}(\beta)\right]$. Note the prefactor $2$, which arises from $\mathcal{O}\left(\frac{1}{N}\right)$ correction to approximating the sum by an integral, is key to the topological correction $\gamma$ in XX spin chain.
@@
$G(\theta)\approx\,\mathrm{e}^{\frac{1}{2}\mathrm{i}\theta N}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-\Re I_{1}(\theta,\beta)\right]}.$ (137)
@@
$G(\pi)$ $\approx 2\cos\left(\frac{N\pi}{2}\right)\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta)-I_{2}(\beta)\right]}.$ (138)
@@ D.2 Away from half filling
where $\nu=\frac{1}{\pi}\arccos\left(-\frac{\mu}{4}\right)$ is the filling fraction.
@@
$\operatorname{Tr}\,\mathrm{e}^{-\beta H}\approx$ $\exp\left[\frac{N}{2\pi}I_{1}(0,\beta,\nu)\right],$ (140)
@@
$I_{1}(\phi,\beta,\nu)$ $=\int_{0}^{2\pi}\ln\left(1+\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{4\beta\cos x-4\beta\cos(\pi\nu)}\right)\mathop{}\!\mathrm{d}x.$ (141)
And for $0<\theta<\pi$, we have
$\left\lvert\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}\right\rvert\approx\exp\left[\frac{N}{2\pi}\Re I_{1}(\theta,\beta,\nu)\right].$ (142)
The phase factor now depends on all the parameters $\theta,\beta$ and $\nu$. For $\theta=\pi$, define
$I_{2}(\beta,\nu)$ $=\int_{0}^{2\pi}\ln\left\lvert 1-\,\mathrm{e}^{4\beta\cos x-4\beta\cos(\pi\nu)}\right\rvert\mathop{}\!\mathrm{d}x.$ (143)
We show in Appendix E that the constant correction relating $I_{2}(\beta,\nu)$ and $\ln\prod\left(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}-4\beta\cos(\pi\nu)}\right)$ is now $2\ln\left[2\sin(\pi\delta)\right]$, where $\delta$ (or $1-\delta$) is the minimal distance between $j$ and the singular point $x=\frac{N\nu}{2}$. For $j$ taking integer values, we can take $\delta=\left[\frac{N\nu}{2}\right]$, which is the decimal part of $\frac{N\nu}{2}$. For $j$ taking half-integer values, we can take $\delta=\left\lvert\left[\frac{N\nu}{2}\right]-\frac{1}{2}\right\rvert$. Thus,
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}$ $\approx-2\sin^{2}\left(\left[\frac{N\nu}{2}\right]\pi\right)\exp\left[I_{2}(\beta,\nu)\right]+2\sin^{2}\left(\left[\frac{N\nu}{2}\right]\pi-\frac{1}{2}\pi\right)\exp\left[I_{2}(\beta,\nu)\right]$ (144)
$=2\cos\left(\nu N\pi\right)\exp\left[\frac{N}{2\pi}I_{2}(\beta,\nu)\right].$ (145)
@@
$\left\lvert G(\theta)\right\rvert\approx\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta,\nu)-\Re I_{1}(\theta,\beta,\nu)\right]}.$ (146)
@@
$G(\pi)\approx 2\cos(\nu N\pi)\,\mathrm{e}^{-\frac{N}{2\pi}\left[I_{1}(0,\beta,\nu)-I_{2}(\beta,\nu)\right]}.$ (147)

### 03 Mar 2024 v2

@@ VI Acknowledgements
Note added: During the finalization of the manuscript, we became aware of closely related works [10] and [44], which study the universal feature in $\alpha(\theta)$. In particular, [10] showed that the O(2) LSM anomaly leads to the cusp at $\theta=\pi$ in $\alpha(\theta)$.
@@ Appendix A Typology of ’t Hooft anomaly in (1+1)d
Next we consider the Levin-Gu model, introducing a $\theta=\pi$ defect [21]:
@@ Appendix C $G(\theta)$ of the Levin-Gu model
We will approximate the discrete products in $Z_{ss^{\prime}}$ by integrals. Define
$I(\phi,\beta)$ $=\int_{0}^{2\pi}\mathop{}\!\mathrm{d}x\,\ln\left(1+\,\mathrm{e}^{\mathrm{i}{\phi}}\,\mathrm{e}^{4\beta\cos x}\right).$ (105)
A crucial property of $I(\phi,\beta)$ for our derivations is that for $-\pi<\phi\leq\pi$,
$\Im I(\phi,\beta)=\pi\phi.$ (106)
@@
$\sum_{j=\frac{s^{\prime}}{2}}^{N-1+\frac{s^{\prime}}{2}}\ln\left(1+\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{\alpha\cos\frac{2\pi j}{N}}\right)=\frac{N}{2\pi}\int_{0}^{2\pi}\mathop{}\!\mathrm{d}x\,\ln(1+\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{\alpha\cos x})+\mathcal{O}(N^{-1}).$ (107)
@@
$Z_{0s^{\prime}}(\phi,\beta)\approx\exp\left[\frac{N}{2\pi}I(\phi,\beta)\right]=\,\mathrm{e}^{\frac{1}{2}\mathrm{i}[\phi]_{\pi}N}\exp\left[\frac{N}{2\pi}\Re I(\phi,\beta)\right].$ (108)
@@
We also need the following inequality between the integrals. Note that
@@
which implies $I(0,\beta)>\Re I\left(\phi,\beta\right)$ for $0<\phi<2\pi$.
@@ C.1 N even
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}$ $\approx\exp\left[\frac{N}{2\pi}I\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]+\exp\left[\frac{N}{2\pi}I\left(-\frac{\theta}{2}+\pi,\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (112)
$\approx\exp\left[\frac{N}{2\pi}I\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (113)
$=\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I\left(\frac{\theta}{2},\beta\right)\right].$ (114)
@@
$Z_{0s^{\prime}}(\pm\pi/2,\beta)=Z_{1s^{\prime}}(\mp\pi/2,\beta)=\exp\left[\frac{N}{2\pi}I\left(\pm\frac{\pi}{2},\beta\right)\right].$ (115)
Crucially, the integrals satisfy $\Re I(\frac{\pi}{2},\beta)=\Re I(-\frac{\pi}{2},\beta)$. Therefore we have
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}\approx 2\cos\left(\frac{\pi N}{4}\right)\exp\left[\frac{N}{2\pi}\Re I(\frac{\pi}{2},\beta)\right].$ (116)
@@
$\approx 4\exp\left[\frac{N}{2\pi}\Re I(\pi,\beta)\right].$ (120)
@@
$G(\theta)\approx\frac{\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I\left(\frac{\theta}{2},\beta\right)\right]}{\exp\left[\frac{N}{2\pi}I(0,\beta)\right]+2(-1)^{\frac{N}{2}}\exp\left[\frac{N}{2\pi}\Re I(\pi,\beta)\right]}\approx\,\mathrm{e}^{\frac{1}{4}\mathrm{i}\theta N}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I(0,\beta)-\Re I\left(\frac{\theta}{2},\beta\right)\right]},$ (121)
@@
$G(\pi)\approx\frac{2\cos\left(\frac{\pi N}{4}\right)\exp\left[\frac{N}{2\pi}\Re I(\frac{\pi}{2},\beta)\right]}{\exp\left[\frac{N}{2\pi}I(0,\beta)\right]+2(-1)^{\frac{N}{2}}\exp\left[\frac{N}{2\pi}\Re I(\pi,\beta)\right]}\approx 2\cos\left(\frac{\pi N}{4}\right)\,\mathrm{e}^{-\frac{N}{2\pi}\left[I(0,\beta)-\Re I(\frac{\pi}{2},\beta)\right]}.$ (122)
@@ C.2 N odd
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}$ $\approx\exp\left[\frac{N}{2\pi}I\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]-\exp\left[\frac{N}{2\pi}I\left(-\frac{\theta}{2}+\pi,\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (124)
$\approx\exp\left[\frac{N}{2\pi}I\left(-\frac{\theta}{2},\beta\right)+\frac{1}{2}\mathrm{i}\theta N\right]$ (125)
$=\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I\left(\frac{\theta}{2},\beta\right)\right],$ (126)
@@
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}\approx$ $\left[\exp\left(\frac{\mathrm{i}\pi N}{4}\right)-\exp\left(\frac{3\mathrm{i}\pi N}{4}\right)\right]\exp\left[\frac{N}{2\pi}\Re I(\frac{\pi}{2},\beta)\right]$ (127)
$=$ $2(-1)^{\frac{N-1}{2}}\sin\left(\frac{\pi N}{4}\right)\exp\left[\frac{N}{2\pi}\Re I(\frac{\pi}{2},\beta)\right].$ (128)
The large $N$ behavior of the partition function is still governed by $I(0,\beta)$ as $I(0,\beta)>\Re I(\pi,\beta)$. Moreover, the two terms $Z_{10}(0,\beta)$ and $Z_{11}(0,\beta)$ have opposite sign for $N$ odd and they cancel each other. Write $N=4m+r$ where $r=1,3$, then $\sin\frac{\pi N}{4}=\sin(m\pi+\frac{r\pi}{4})=(-1)^{m}\sin\frac{r\pi}{4}=\frac{1}{\sqrt{2}}(-1)^{m}$. We conclude that for $0<\theta<\pi$,
$G(\theta)\approx\frac{\exp\left(\frac{1}{4}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I\left(\frac{\theta}{2},\beta\right)\right]}{\exp\left[\frac{N}{2\pi}I(0,\beta)\right]}=\,\mathrm{e}^{\frac{1}{4}\mathrm{i}\theta N}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I(0,\beta)-\Re I\left(\frac{\theta}{2},\beta\right)\right]},$ (129)
@@
$G(\pi)\approx\frac{(-1)^{m+\frac{r-1}{2}}\sqrt{2}\exp\left[\frac{N}{2\pi}\Re I(\frac{\pi}{2},\beta)\right]}{\exp\left[\frac{N}{2\pi}I(0,\beta)\right]}=(-1)^{m+\frac{r-1}{2}}\sqrt{2}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I(0,\beta)-\Re I(\frac{\pi}{2},\beta)\right]}.$ (130)
@@ D.1 Half filling
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}\approx\exp\left[\frac{N}{2\pi}I\left(\theta,\beta\right)\right]=\exp\left(\frac{1}{2}\mathrm{i}\theta N\right)\exp\left[\frac{N}{2\pi}\Re I\left(\theta,\beta\right)\right].$ (132)
@@
We have proved that $\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}=0$ for $N$ odd, so we only need to consider $N$ even case. $Z_{00}(0,\beta)$ and $Z_{01}(0,\beta)$ can be approximated by Eq. (108), and they cancel each other. The other two terms are already considered in Appendix C, and we have $\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}=2(-1)^{\frac{N}{2}}\exp\left[\frac{N}{2\pi}\Re I(\pi,\beta)\right]$. Note the prefactor $2$, which arises from $\mathcal{O}\left(\frac{1}{N}\right)$ correction to approximating the sum by an integral, is key to the topological correction $\gamma$ in XX spin chain.
@@
$G(\theta)\approx\,\mathrm{e}^{\frac{1}{2}\mathrm{i}\theta N}\,\mathrm{e}^{-\frac{N}{2\pi}\left[I(0,\beta)-\Re I(\theta,\beta)\right]}.$ (134)
@@
$G(\pi)$ $\approx 2\cos\left(\frac{N\pi}{2}\right)\,\mathrm{e}^{-\frac{N}{2\pi}\left[I(0,\beta)-\Re I(\pi,\beta)\right]}.$ (135)
@@ D.2 Away from half filling
where $\nu=\frac{1}{\pi}\arccos\left(-\frac{\mu}{4}\right)$ is the filling fraction at zero temperature.
@@
$\operatorname{Tr}\,\mathrm{e}^{-\beta H}\approx$ $\exp\left[\frac{N}{2\pi}I(0,\beta,\nu)\right],$ (137)
@@
$I(\phi,\beta,\nu)$ $=\int_{0}^{2\pi}\ln\left(1+\,\mathrm{e}^{\mathrm{i}\phi}\,\mathrm{e}^{4\beta\cos x-4\beta\cos(\pi\nu)}\right)\mathop{}\!\mathrm{d}x.$ (138)
For $0<\theta<\pi$, we have
$\left\lvert\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\theta Q}\,\mathrm{e}^{-\beta H}\right\rvert\approx\exp\left[\frac{N}{2\pi}\Re I(\theta,\beta,\nu)\right].$ (139)
The phase factor now depends on all the parameters $\theta,\beta$ and $\nu$. For $\theta=\pi$, we show in Appendix E that the constant correction relating $I(\pi,\beta,\nu)$ and $\ln\prod\left(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}-4\beta\cos(\pi\nu)}\right)$ is now $2\ln\left[2\sin(\pi\delta)\right]$, where $\delta$ (or $1-\delta$) is the minimal distance between $j$ and the singular point $x=\frac{N\nu}{2}$. For $j$ taking integer values, we can take $\delta=\left[\frac{N\nu}{2}\right]$, which is the decimal part of $\frac{N\nu}{2}$. For $j$ taking half-integer values, we can take $\delta=\left\lvert\left[\frac{N\nu}{2}\right]-\frac{1}{2}\right\rvert$. Thus,
$\operatorname{Tr}\,\mathrm{e}^{\mathrm{i}\pi Q}\,\mathrm{e}^{-\beta H}$ $\approx-2\sin^{2}\left(\left[\frac{N\nu}{2}\right]\pi\right)\exp\left[\Re I(\pi,\beta,\nu)\right]+2\sin^{2}\left(\left[\frac{N\nu}{2}\right]\pi-\frac{1}{2}\pi\right)\exp\left[\Re I(\pi,\beta,\nu)\right]$ (140)
$=2\cos\left(\nu N\pi\right)\exp\left[\frac{N}{2\pi}\Re I(\pi,\beta,\nu)\right].$ (141)
@@
$\left\lvert G(\theta)\right\rvert\approx\,\mathrm{e}^{-\frac{N}{2\pi}\left[I(0,\beta,\nu)-\Re I(\theta,\beta,\nu)\right]}.$ (142)
@@
$G(\pi)\approx 2\cos(\nu N\pi)\,\mathrm{e}^{-\frac{N}{2\pi}\left[I(0,\beta,\nu)-\Re I(\pi,\beta,\nu)\right]}.$ (143)